OpenStudy (anonymous):

Solve the inequality \[x^4 ge x^2\] and express the solution as an interval.

OpenStudy (anonymous):

\[x^4 \ge x^2\] \[x^4 - x^2\ge 0\]

OpenStudy (anonymous):

x^4 is greater than or equal to x^2? x is a function of all real numbers.... (-inf,inf)

OpenStudy (anonymous):

wait....I'm sorry...I messed up.

OpenStudy (anonymous):

x^4 is greater than or equal to x^2 I believe this means that x has to be greater than or equal to 1

OpenStudy (anonymous):

For example.... (1/2)^4=1/16 (1/2)^2=1/4

OpenStudy (anonymous):

For the interval I got [-1,0) U [1,infinity)

OpenStudy (anonymous):

it would be (-inf,-1] U [1, inf)

OpenStudy (anonymous):

 can you take it from here?